OpenStudy (anonymous):

which expression represents x^2-3x-10 over x^2-25

OpenStudy (anonymous):

Options please

OpenStudy (anonymous):

x-5 is the answer i think!

OpenStudy (anonymous):

that is the only expression that is in common for both the expressions!

OpenStudy (anonymous):

(x^2-3x-10)=(x-5)(x+2) (x^2-25)=(x-5)(x-5) So divide them and get your answer

OpenStudy (anonymous):

or x-2 over x-5

OpenStudy (anonymous):

x^2-3x-10/x^2-25 =(x^2+2x-5x-10)/(x-5)(x+5) =x(x+2)-5(x+2)/(X-5)(X+5) =(x-5)(x+2)/(x-5)(x+5) =(x+2)/(x+5)

OpenStudy (anonymous):

its x-5

OpenStudy (anonymous):

no it's (x+2)/(x+5)

OpenStudy (anonymous):

i got  but i ot sure

OpenStudy (anonymous):

yeah that's the answer

OpenStudy (anonymous):

ok i thought u are asking for expression in comon!

OpenStudy (anonymous):

but u said it was (x+2)/(x+5)

OpenStudy (anonymous):

oh wait let me check

OpenStudy (anonymous):

yeah wait it is. you got yours almost right. just you're suppose to have positives at the end. check my steps

OpenStudy (anonymous):

thanks